# Generative Inverse Design of Metamaterials with Functional Responses by Interpretable Learning

### v1

@@ 3.2 Applying RIGID to Design Optical Metasurfaces with Target High-Absorbance Wavelengths
In this case study, we consider four types of cross-sections ($c\in\{1,2,3,4\}$) chosen from the literature (Fig. 4B). It is assumed that a 3D geometric instance is composed of a stack of three layers of prismatic unit cells, each of which is vertically extruded and stacked (Fig. 4A). The geometries constructed in this way can be regarded as an instantiation of multilayered metasurfaces [52, 53, 54, 55, 56], which offer richer design freedom than the single-layer counterpart. The height of each layer ($h_{l},l=1,2,3$) is allowed to continuously vary between 50 and 150 nm. Herein we do not consider parametric variations of a given type of unit cell cross-section; yet those can be trivially incorporated in the proposed design framework if necessary.
@@
Figure 4: Optical metasurface design problem configuration and results. (A-B) Design variables (materials, layer heights, and cross-section geometry types). (C) A sample absorbance spectrum and the wavelength intervals (highlighted wavelength regions) corresponding to absorbance above the threshold $t$. The design objective is to generate new optical metasurface designs that exhibit higher absorbance than a threshold $t$ at the user-defined wavelength interval(s). (D) KDE of the estimated likelihood for generated designs. (E) Validation metrics under varying sampling thresholds. (F) Designs (geometries and material selections) and corresponding absorbance spectra of five generated metasurfaces with the highest likelihood of satisfying specified target high-absorbance regions (marked as highlighted wavelength regions). Here all five designs satisfy the target. Generated designs for other targets are shown in Appendix, Figs. 11-12.
@@
We also design the material of each layer ($m_{l},l=1,2,3$). Three dielectric materials of interest, each of which is assigned to a different color in Fig. 4A, are Ti (red), Si (blue), and Ge (yellow). In general, a dielectric material is characterized through a complex refractive index $\tilde{n}\in\mathbb{C}$ defined as $\tilde{n}=n+jk$, where $j=\sqrt{-1}$ is the imaginary unit, $n\in\mathbb{R}$ involves the speed at which the light propagates through the material, and $k\in\mathbb{R}$ is the extinction coefficient that dictates the energy loss due to the material. Within the frequency regime of interest, those exhibit nonlinear dispersion; both the real and imaginary terms in general are a non-analytic function of excitation wavelength $s$, i.e., $n(s)$ and $k(s)$. In addition, the impact of the same material choice on the spectral response $A(s)$ varies depending on the layer location at which the material is placed. Thus the highlight of this case study is the combinatorial nature of design, whose spectral responses are affected by the joint contributions of geometry and material.
@@
Based on the above configuration, we randomly sampled 258 sets of design variables $\mathbf{x}=(c,h_{1},h_{2},h_{3},m_{1},m_{2},m_{3})$ and computed their corresponding absorbance spectra using wave analysis. We set $t=0.9$ as the absorbance threshold, so that “high” absorbance means the absorbance $A(s)$ is no less than 0.9. We trained a random forest on 206 training data (i.e., 80% of the 258 designs and corresponding absorbance spectra) to predict whether “high” absorbance is presented (i.e., the binary indicator $y=1$) at a wavelength $s$ for a given design $\mathbf{x}$. The trained random forest has a test F1 score of 0.83. The confusion matrix on test data is shown in Appendix, Tab. 2.
@@
Figure 4F shows generated optical metasurface designs with the top five likelihood estimations for a randomly created target. While the materials, cross-section geometries, and layer heights of generated designs can be different, all the designs satisfy the target (Fig. 4F, right panel). To further enhance the diversity of final solutions, we can use sampling strategies such as the one proposed in Ref. [60, 61] to identify a subset (of generated solutions) that simultaneously exhibits high likelihood and high diversity. Generated designs based on the other nine targets can be found in Appendix, Figs. 11-12.
@@ 3.3 Synthetic Design Problems for Rapid Validation and Visualization
where $z$ represents quantitative response and $t$ is a threshold that converts $z$ into a qualitative response $I(a,b;s)$. Specifically, $I(a,b;s)=1$ can mean the existence of a bandgap or high absorbance at a frequency $s$. Then $\{s|I(a,b;s)=1\}$ represents a range of $s$ that mimics our design targets, such as the bandgap or the frequency range of high absorbance. By varying $a$ and $b$, we can produce different synthetic responses and ranges. Therefore, we can use $a$ and $b$ as synthetic design variables. There is a clear relation between these design variables and the range that Eq. 2 creates — $a$ and $b$ control the center location and the width of the range, respectively.
@@
For each synthetic example, we split the data into 80 training data and 20 test data, and trained a random forest, with the same hyperparameter settings as the other problems, to predict the binary indicators $I(a,b;s)$. The F1 scores are 0.85 and 0.86 for the SqExp and the SupSin problems, respectively. The resulting confusion matrices are shown in Appendix, Tables 3-4. We evaluated the inverse design performance with the trained models.
Figure 6: Visualization of estimated likelihood and validation metrics for synthetic problems. (A) Likelihood function values for randomly created design targets. Orange lines show boundaries of actual feasible regions associated with the targets. (B) Likelihood function values estimated by random forests with varying numbers of decision trees. The design target is set as $\{[0.45,0.48]\}$ for the SqExp problem and $\{[0.63,0.68],[0.69,0.71]\}$ for the SupSin problem. (C) Validation metrics for inverse design generation.
Due to the fast evaluation of Equations 2 and 3, we can exhaust all the possible solutions in the design space to obtain the ground-truth feasible region(s) for a target. Figure 6A shows the estimated likelihood values and the ground-truth feasible regions under randomly created targets. In general, high-likelihood regions match actual feasible regions well, which further demonstrates the effectiveness of RIGID. We can also observe that feasible regions in the SqExp and the SupSin problems follow distinct patterns. In the SqExp problem, $a$ and $b$ control the center location and the width of the output range, respectively. Therefore, the position of the feasible region along the $a$-axis moves with the location of the target range, while the feasible region gradually shrinks as $b$ decreases since the decrease of $b$ (i.e., output range width) causes the choice of $a$ (i.e., output range center location) to be more restricted to fit the target range. In the SupSin problem, there might be multiple bandgaps appearing at the peaks of the superposed sine function in Eq. 3. Design variables $a$ and $b$ control bandgap locations by translating each sine function. Due to the sine function’s periodicity, we can obtain multiple feasible regions along both $a$- and $b$-axes. Figure 6A shows that the likelihood estimation by RIGID successfully captured the above-mentioned patterns of feasible regions.
@@ 4 Summary and Outlook
We proposed RIGID, a single-shot inverse design method that generates metamaterials to satisfy qualitative behaviors of functional responses. Such qualitative behaviors are important design targets in many applications such as tailoring bandgaps of acoustic metamaterials for wave-guiding, focusing, and vibration mitigation, or tailoring the absorption level of optical metasurfaces at certain wavelengths for medical sensing, imaging, and communication applications. Unlike most existing inverse design methods that require training an inverse model to map targets to designs, the RIGID method takes advantage of the random forest’s interpretability and derives the likelihood of target satisfaction by probing the trained forward model. Incorporated with MCMC, one can sample a desired number of new designs based on the estimated likelihood. Therefore, RIGID functions as a generative model that can capture the conditional distribution of satisfying designs given a target, or in other words, the one-to-many mapping from the target to satisfying designs. Using both real-world and synthetic design problems, we demonstrated that RIGID is efficient and effective on datasets with training sample sizes smaller than 250. Thus, RIGID is particularly useful when data collection is expensive, as in many cases where high-fidelity simulation or experimental data are needed. We used both qualitative and quantitative approaches to validate the proposed method. The quantitative results revealed a strong correlation between the estimated likelihood of a solution and its probability of actually achieving the target, which demonstrated the effectiveness of the likelihood estimation. Due to the fast evaluation of output responses and the transparency of ground-truth solutions, the proposed synthetic problems can be used for future benchmarking studies of metamaterial design problems.
@@
While we address qualitative design targets in this study, the idea of using random forest-based models for inverse design has the potential to generalize to quantitative targets. Such problems can be, for example, generating optical metasurface designs with specific optical spectra [62, 21], generating functional materials with target nonlinear constitutive relations [33, 63], or generating programmable metamaterials with prescribed functional responses [64, 65]. It is also straightforward to adjust the target to achieve multifunctionality (e.g., negative/positive Poisson’s ratio under low/high compression rate [66]).
@@ Acknowledgement
Appendix A Related Works on Single-Shot Inverse Design
Figure 7 shows the schematic diagrams of three mainstream machine learning models for single-shot inverse design of metamaterials.
Figure 7: Three major machine learning models for single-shot inverse design. Note that these are only conceptual plots and the networks can be more complicated in reality. Also, due to the diversity and complexity of conditional generative model architectures, we only show the network for design generation based on commonly used models (e.g., generators in cGANs or decoders in cVAEs), while omitting networks that are necessary for training (e.g., discriminators in cGANs or encoders in cVAEs).
Appendix B Explanation of Step 2 in RIGID
@@
Appendix C Computational Methods for Data Collection and Validation
C.1 Computation of Acoustic Dispersion Relations
[…]
C.2 Computation of Optical Absorbance Spectra
[…]
Appendix D Test performance of forward prediction
@@
In the acoustic metamaterial design problem, we have 57 test designs, yielding 5,700 test points as the entire frequency range of dispersion relations is discretized into 100 intervals for each design. In the optical metasurface design problem, we have 52 test designs, with each functional response discretized into 33 points, which yields 1,716 test points in total. In each of the synthetic design problems, we have 20 synthetic test designs, with each synthetic response discretized into 100 points. This results in 2,000 test data points in total. Confusion matrices showing test performances are in Tables 1-4.
Table 1: Confusion matrix on test data in the acoustic metamaterial design problem. The test F1 score is 0.82.
@@
Table 2: Confusion matrix on test data in the optical metasurface design problem. The test F1 score is 0.83.
@@
Table 3: Confusion matrix on test data in the SqExp problem. The test F1 score is 0.85.
@@
Table 4: Confusion matrix on test data in the SupSin problem. The test F1 score is 0.86.
@@
Appendix E Generated Designs and Corresponding Responses
Figures 8-12 show the rest of generated designs and their corresponding responses in addition to those in Figures 3F and 4F.
@@ Appendix E Generated Designs and Corresponding Responses
Figure 8: Geometries and corresponding dispersion relations of generated acoustic metamaterial designs for bandgap targets \@slowromancapii@-\@slowromancapiv@ (marked as gray-shaded frequency regions). Each row shows five designs with the highest likelihood of satisfying a target.
@@
Figure 9: Geometries and corresponding dispersion relations of generated acoustic metamaterial designs for bandgap targets \@slowromancapv@-\@slowromancapvii@ (marked as gray-shaded frequency regions). Each row shows five designs with the highest likelihood of satisfying a target.
@@
Figure 10: Geometries and corresponding dispersion relations of generated acoustic metamaterial designs for bandgap targets \@slowromancapviii@-\@slowromancapx@ (marked as gray-shaded frequency regions). Each row shows five designs with the highest likelihood of satisfying a target.
@@
Figure 11: Designs and corresponding absorbance spectra of generated optical metasurface designs for design targets \@slowromancapii@-\@slowromancapvi@ (marked as gray-shaded wavelength regions). Each row shows five designs with the highest likelihood of satisfying a target (left panel) and their absorbance spectra (right panel).

### v2

@@ 3.2 Applying RIGID to Design Optical Metasurfaces with Target High-Absorbance Wavelengths
Figure 4: Optical metasurface design problem configuration and results. (A-B) Design variables (materials, layer thicknesses, and cross-section geometry types). (C) A sample absorbance spectrum and the wavelength intervals (highlighted wavelength regions) corresponding to absorbance above the threshold $t$. The design objective is to generate new optical metasurface designs that exhibit higher absorbance than a threshold $t$ at the user-defined wavelength interval(s). (D) KDE of the estimated likelihood for generated designs. (E) Validation metrics under varying sampling thresholds. (F) Designs (geometries and material selections) and corresponding absorbance spectra of five generated metasurfaces with the highest likelihood of satisfying specified target high-absorbance regions (marked as highlighted wavelength regions). All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm. Here all five designs satisfy the target. Generated designs for other targets are shown in SI Appendix, Figs. S7-S9.
In this case study, we consider four types of cross-sections ($c\in\{1,2,3,4\}$) chosen from the literature (Fig. 4B). It is assumed that a 3D geometric instance is composed of a stack of three layers of prismatic unit cells, each of which is vertically extruded and stacked (Fig. 4A). The geometries constructed in this way can be regarded as an instantiation of multilayered metasurfaces [68, 69, 70, 71, 72], which offer richer design freedom than the single-layer counterpart. The thickness of each layer ($h_{l},l=1,2,3$) is allowed to continuously vary between 50 and 150 nm. Herein we do not consider parametric variations of a given type of unit cell cross-section; yet those can be trivially incorporated in the proposed design framework if necessary.
@@
We also design the material of each layer ($m_{l},l=1,2,3$). Three dielectric materials of interest, each of which is assigned to a different color in Fig. 4A, are Ti (red), Si (blue), and Ge (yellow). The challenge associated with this design problem is its mixed-variable design space containing three continuous variables (i.e., layer thicknesses) and four categorical variables (i.e., material choices at each layer and the geometry type), which may lead to potentially complicated underlying design-response relations (e.g., those across different geometry or material types). In general, a dielectric material is characterized through a complex refractive index $\tilde{n}\in\mathbb{C}$ defined as $\tilde{n}=n+jk$, where $j=\sqrt{-1}$ is the imaginary unit, $n\in\mathbb{R}$ involves the speed at which the light propagates through the material, and $k\in\mathbb{R}$ is the extinction coefficient that dictates the energy loss due to the material. Within the frequency regime of interest, those exhibit nonlinear dispersion; both the real and imaginary terms in general are a non-analytic function of excitation wavelength $s$, i.e., $n(s)$ and $k(s)$. In addition, the impact of the same material choice on the spectral response $A(s)$ varies depending on the layer location at which the material is placed.
@@
Based on the above configuration, we randomly sampled 258 sets of design variables $\mathbf{x}=(c,h_{1},h_{2},h_{3},m_{1},m_{2},m_{3})$ and computed their corresponding absorbance spectra using wave analysis. We set $t=0.9$ as the absorbance threshold, so that “high” absorbance means the absorbance $A(s)$ is no less than 0.9. We trained a random forest on 206 training data (i.e., 80% of the 258 designs and corresponding absorbance spectra) to predict whether “high” absorbance is presented (i.e., the binary indicator $y=1$) at a wavelength $s$ for a given design $\mathbf{x}$. The trained random forest has a test F1 score of 0.83. The confusion matrix on test data is shown in SI Appendix, Tab. S2.
@@
Figure 4F shows generated optical metasurface designs with the top five likelihood estimations for a randomly created target. While the materials, cross-section geometries, and layer thicknesses of generated designs can be different, all the designs satisfy the target (Fig. 4F, right panel). To further enhance the diversity of final solutions, we can use sampling strategies such as the one proposed in Ref. [76, 77] to identify a subset (of generated solutions) that simultaneously exhibits high likelihood and high diversity. Generated designs based on the other nine targets can be found in SI Appendix, Figs. S7-S9.
@@ 3.3 Synthetic Design Problems for Rapid Validation and Visualization
where $z$ represents quantitative response and $t$ is a threshold that converts $z$ into a qualitative response $I(a,b;s)$. Specifically, $I(a,b;s)=1$ can mean the existence of a bandgap or high absorbance at a frequency $s$. Then $\{s|I(a,b;s)=1\}$ represents a range of $s$ that mimics our interested material behavior, such as the frequency range of the bandgap or the wavelength range of high absorbance. We call this the positive range. By varying $a$ and $b$, we can produce different synthetic responses and positive ranges. Therefore, we can use $a$ and $b$ as synthetic design variables. There is a clear relation between these design variables and the positive range that Eq. 2 creates — $a$ and $b$ control the center location and the width of the positive range, respectively.
@@
Figure 6: Visualization of estimated likelihood and validation metrics for synthetic problems. (A) Likelihood function values for randomly created design targets. Orange lines show boundaries of actual feasible regions associated with the targets $\mathcal{T}=\{I(a,b;s)=1|\forall s\in\Omega\}$. (B) Likelihood function values estimated by random forests with varying numbers of decision trees. The design target is set as $\mathcal{T}=\{I(a,b;s)=1|\forall s\in[0.45,0.48]\}$ for the SqExp problem and $\mathcal{T}=\{I(a,b;s)=1|\forall s\in[0.63,0.68]\cup[0.69,0.71]\}$ for the SupSin problem. (C) Validation metrics for inverse design generation.
For each synthetic example, we split the data into 80 training data and 20 test data, and trained a random forest, with the same hyperparameter settings as the other problems, to predict the binary indicators $I(a,b;s)$. The F1 scores are 0.85 and 0.86 for the SqExp and the SupSin problems, respectively. The resulting confusion matrices are shown in SI Appendix, Tables S3-S4. We evaluated the inverse design performance with the trained models.
Due to the fast evaluation of Equations 2 and 3, we can exhaust all the possible solutions in the design space to obtain the ground-truth feasible region(s) for a target. Figure 6A shows the estimated likelihood values and the ground-truth feasible regions under randomly created targets. In general, high-likelihood regions match actual feasible regions well, which further demonstrates the effectiveness of RIGID. We can also observe that feasible regions in the SqExp and the SupSin problems follow distinct patterns. In the SqExp problem, $a$ and $b$ control the center location and the width of the positive range, respectively. Therefore, the position of the feasible region along the $a$-axis moves with the location of the target range, while the feasible region gradually shrinks as $b$ decreases since the decrease of $b$ (i.e., positive range width) causes the choice of $a$ (i.e., positive range center location) to be more restricted to fit the target range. In the SupSin problem, there might be multiple positive ranges appearing at the peaks of the superposed sine function in Eq. 3. Design variables $a$ and $b$ control positive range locations by translating each sine function. Due to the sine function’s periodicity, we can obtain multiple feasible regions along both $a$- and $b$-axes. Figure 6A shows that the likelihood estimation by RIGID successfully captured the above-mentioned patterns of feasible regions.
@@ 4 Summary and Outlook
We proposed RIGID, an iteration-free, single-shot inverse design method that generates metamaterials to satisfy qualitative behaviors of functional responses. Such qualitative behaviors are important design targets in many applications such as tailoring bandgaps of acoustic metamaterials for wave-guiding, focusing, and vibration mitigation, or tailoring the absorption level of optical metasurfaces at certain wavelengths for medical sensing, imaging, and communication applications. Unlike most existing machine learning-based inverse design methods that require training an inverse model to map targets to designs, the RIGID method takes advantage of the random forest’s interpretability and derives the likelihood of target satisfaction by probing the trained forward model. Incorporated with MCMC, one can sample a desired number of new designs based on the estimated likelihood. Therefore, RIGID functions as a generative model that can capture the conditional distribution of satisfying designs given a target, or in other words, the one-to-many mapping from the target to satisfying designs. Using both real-world and synthetic design problems, we demonstrated that RIGID is efficient and effective on datasets with training sample sizes smaller than 250. We used both qualitative and quantitative approaches to validate the proposed method. The quantitative results revealed a strong correlation between the estimated likelihood of a solution and its probability of actually achieving the target, which demonstrated the effectiveness of the likelihood estimation. Due to the fast evaluation of output responses and the transparency of ground-truth solutions, the proposed synthetic problems can be used for future benchmarking studies of metamaterial design problems.
RIGID offers two levels of interpretability. Firstly, RIGID is more transparent than deep learning models. The transparency of the random forest as a forward model helps us understand how the predictions are made. The inverse design process based on likelihood estimation and sampling is also transparent. The transparency of both forward and inverse processes offers insights into the decision-making process and diagnostic capabilities. Secondly, owing to the interpretability of the forward prediction model, we can derive the likelihood, which the inverse design relies on, by probing into each decision path of trees in the random forest. The likelihood is an interpretable metric to help us understand the probability of a generated design satisfying the given target and thus aid the decision-making process when selecting final solutions. The results (Figures 3D, 3E, 4D, 4E, and 6C) have demonstrated a high correlation between the likelihood and actual probability of target satisfaction. The likelihood also serves as an indicator of the need for adjusting the design space. When the likelihood is low across the entire design space, the model believes the target is almost impossible to achieve, which may indicate a need for expanding the design space (i.e., extending design space bounds or adding new design freedom) to allow wider coverage of the response space and easier discovery of satisfying solutions. Please see more discussion in SI Appendix, Sec. D. Another benefit of having the explicit likelihood as an interpretable metric is that we can adjust the likelihood threshold to trade-off exploration and exploitation — higher (lower) threshold means more focus on exploitation (exploration). Both levels of interpretability were absent from most prior works of iteration-free, single-shot inverse design methods for metamaterials.
RIGID is particularly useful when data collection is expensive, as in many cases where high-fidelity simulation or experimental data are needed. Although hardware-accelerated simulation and surrogate solvers have advanced in recent years to make simulation faster without sacrificing too much accuracy, the trade-off between time/computational resources and accuracy will always pose a problem. For example, hardware acceleration for high-fidelity simulation relies on high-end GPUs/CPUs and other high-performance computing infrastructure, which might not be accessible under limited budgets. With surrogate solvers, the need for computational resources reduces, but so will the accuracy. If the surrogate solver is data-driven, we will still have the problem of collecting high-fidelity data in the first place. Thus, we believe that the issue of data deficiency will always exist, especially when high-fidelity data is desired but time and computational resources are constrained. With most of the prior inverse design works based on deep learning with a high demand for data and computational resources, it is useful to create a method that works under scenarios at the other end of the data/computational resource requirement spectrum. Additionally, our proposed method is not only useful with high-fidelity simulation/numerical data, but also with experimental data. When it is difficult to build an accurate numerical model, the only way to acquire high-fidelity data is through experiments. In this case, data deficiency will become a pressing challenge for deep learning-based methods.
This work focuses on the scenario of small data, low-dimensional problems. As the problem dimension increases, the requirement for dataset size will quickly increase due to the curse of dimensionality. While there is a large collection of deep learning methods to address high-dimensional and large dataset problems, the setting of small data and low dimensionality seems to be under-studied by prior works, despite this setting’s practicality under constrained budgets. RIGID can fill this gap.
The MCMC in RIGID can sample new designs in the entire design space based on the estimated likelihood, which is not limited to local optimality and has the potential to discover multiple very different satisfying solutions (if any) in one shot. This is clearly demonstrated in the SupSin example where there are multiple disconnected feasible regions in the design space and the likelihood estimated by RIGID successfully coincides with these feasible regions (Fig. 6A). The ability to generate multiple satisfying solutions allows us to consider other figures of merit, such as manufacturability and cost, in addition to the main objective of satisfying the target response. We can also control the likelihood threshold to adjust the diversity of generated designs.
The use of a random forest in RIGID allows easy consideration of discrete design variables. For example, our optical metasurface design problem has the geometry type and the material choice for each layer as categorical design variables. In contrast, these discrete variables will pose challenges for other inverse design methods based on some machine learning models (e.g., conditional generative models) or gradient-based optimization. On the other hand, the simplicity of the tree-based model does not hurt its capability of solving complex problems, as a decision tree can represent arbitrarily complex decision boundaries.
@@
While we address qualitative design targets in this study, a similar idea has the potential to generalize to quantitative targets. Such problems can be, for example, generating optical metasurface designs with specific optical spectra [78, 23], generating functional materials with target nonlinear constitutive relations [41, 79], or generating programmable metamaterials with prescribed functional responses [80, 81]. It is also straightforward to adjust the target to achieve multifunctionality (e.g., negative/positive Poisson’s ratio under low/high compression rate [82]).
@@
5 Experimental Section
Computation of Acoustic Dispersion Relations
We performed Bloch-wave analysis in COMSOL Multiphysics to compute the dispersion relations of acoustic metamaterials. Poisson’s ratio of 0.49, Young’s modulus of 2.7 GPa, and density of 1170 kg/m3 were set as material properties with $\sim 1.5\times 10^{4}$ mesh elements per unit cell. We used Floquet-Bloch periodic boundary conditions to obtain the first 60 eigenfrequencies along all symmetry domains of the cubic irreducible Brillouin zone (Fig. 3B) for all lattices, thus generating a dispersion relation.
Computation of Optical Absorbance Spectra
We computed the absorbance spectra for optical metasurfaces using wave analysis inspired by Zhang et al. [72]. The RF Module of COMSOL Multiphysics® [86] was used to evaluate the spectral response of concern, which is the energy absorbance $A(s)$ in the visible regime (380-700 nm). An absorbance spectrum is computed with respect to 33 wavelength components $s_{k}$ that are uniformly discretized over the specified range. An incident plane wave is assumed to be given from the port, located at the top face of the analysis domain. We set the periodicity of the analysis domain as 400 nm. The periodic boundary condition on electromagnetic fields is imposed on the lateral faces of the analysis domain. A substrate made of SiO2 is placed right below a given unit cell instance (the black layers in Fig. 4A-B). With full electric fields computed through the wave analysis, the energy absorbance at a single wavelength $s$ is quantified as $A(s)=1-|S_{11}(s)|^{2}$, where $S_{ij}$ is the component of the $S$-parameter matrix that specifies energy transfer between ports. We use the data presented in Ref. [87] to set the material dispersion of the dielectric.
@@ Acknowledgement
Data Availability Statement
Source code and data for this work are available at https://github.com/DIGITLab23/RIGID.
Keywords
metamaterials, inverse design, generative design, random forest, interpretable machine learning, functional response
Supporting Information for
Generative Inverse Design of Metamaterials with Functional Responses by Interpretable Learning
Appendix S1 Existing Works on Deep Learning-Based Iteration-Free, Single-Shot Inverse Design
The three mainstream deep learning methods for iteration-free, single-shot metamaterial inverse design are direct inverse mapping, Tandem Neural Network (T-NN), and conditional generative models. Most prior works fall into one of the three categories or at the intersection of two categories. Here we give more details of each method. The direct inverse mapping approach (Fig. S1A) simply learns a one-to-one mapping from responses to designs. It fails to account for non-unique solutions of the inverse problem, which leads to conflicting labels (i.e., designs) for the same input (i.e., response) and hence causes training convergence issues.
The T-NN (Fig. S1B) was designed to solve this non-uniqueness issue. Its training is split into two steps: (1) pretraining the forward-modeling network to approximate the design-response mapping and (2) training the cascaded network by freezing the weights of the pretrained forward-modeling network. There is no loss function that forces designs at the intermediate layer to match data (which contains conflicting instances), hence the training convergence issue is avoided. However, the original T-NNs still learn a one-to-one response-design mapping and cannot generate multiple satisfying solutions. Variants of T-NN that learn one-to-many mappings are mentioned in the main text.
The conditional generative models (Fig. S1C) are the most common approach to learning a one-to-many mapping from responses to designs. Different generative models have distinct ways of learning conditional distributions. In general, this is realized by training neural networks to transform responses and random noise (or latent variables) into designs, so that the trained network can generate a non-deterministic design solution from a given target response and randomly sampled noise, which is equivalent to sampling from a conditional distribution.
Figure S1: Three major machine learning models for single-shot inverse design. Note that these are only conceptual plots and the networks can be more complicated in reality. Also, due to the diversity and complexity of conditional generative model architectures, we only show the network for design generation based on commonly used models (e.g., generators in cGANs or decoders in cVAEs), while omitting networks that are necessary for training (e.g., discriminators in cGANs or encoders in cVAEs).
Appendix S2 Explanation of Step 2 in RIGID
@@
Appendix S3 Test Performance of Forward Prediction
@@
In the acoustic metamaterial design problem, we have 57 test designs, yielding 5,700 test points as the entire frequency range of dispersion relations is discretized into 100 intervals for each design. In the optical metasurface design problem, we have 52 test designs, with each functional response discretized into 33 points, which yields 1,716 test points in total. In each of the synthetic design problems, we have 20 synthetic test designs, with each synthetic response discretized into 100 points. This results in 2,000 test data points in total. Confusion matrices showing test performances are in Tables S1-S4.
Table S1: Confusion matrix on test data in the acoustic metamaterial design problem. The test F1 score is 0.82.
@@
Table S2: Confusion matrix on test data in the optical metasurface design problem. The test F1 score is 0.83.
@@
Table S3: Confusion matrix on test data in the SqExp problem. The test F1 score is 0.85.
@@
Table S4: Confusion matrix on test data in the SupSin problem. The test F1 score is 0.86.
@@
Appendix S4 Generated Designs and Corresponding Responses
In each of the acoustic and optical metamaterial design problems, we randomly created 10 design targets (i.e., frequency ranges within which bandgaps exist or wavelength ranges within which absorption is high). Given each design target, we generate multiple design solutions but only visualize five solutions having the highest likelihood values. The main text shows the five solutions for the first design target in each problem. Figures S2-S9 show the top solutions (ranked by likelihood) for the other design targets.
While these limited examples cannot reveal the same degree of generalizable insights as the satisfaction rate and the average score do, they still offer an intuitive glimpse into the method’s performance and functionality. Most of these generated designs satisfy their corresponding targets. Some distinct designs can achieve the same target, demonstrating the capability of RIGID in solving the non-uniqueness issue of inverse design. For some targets, even the maximum likelihood of generated designs is extremely small (e.g., Target \@slowromancapiv@ shown in the first row in Fig. S4), indicating small likelihood values across the entire design space. This can be owing to the difficulty of achieving the target under the limit of the design space. For example, all the top five generated designs for Target \@slowromancapiv@ have a likelihood of 0.001. This means that the model believes satisfying the target is almost impossible, which may be due to the wide span of the combination of two target bandgaps. This difficulty may result from the limitation of the current design space. A larger design space can lead to a wider coverage of the response space and hence more likely to contain satisfactory solutions. Based on the estimated likelihood, we can decide whether expanding the design space (i.e., extending design space bounds or adding new design freedom) is necessary to discover satisfying solutions.
@@
Figure S2: Geometries and corresponding dispersion relations of generated acoustic metamaterial designs for bandgap targets \@slowromancapii@ and \@slowromancapiii@ (marked as gray-shaded frequency regions). Each row shows five designs with the highest likelihood of satisfying a target. All the radii ($r_{\text{strut}}$, $r_{\text{center}}$, and $r_{\text{corner}}$) have a unit of $\mu$m.
@@
Figure S3: Geometries and corresponding dispersion relations of generated acoustic metamaterial designs for bandgap targets \@slowromancapiv@ and \@slowromancapv@ (marked as gray-shaded frequency regions). Each row shows five designs with the highest likelihood of satisfying a target. All the radii ($r_{\text{strut}}$, $r_{\text{center}}$, and $r_{\text{corner}}$) have a unit of $\mu$m.
@@
Figure S4: Geometries and corresponding dispersion relations of generated acoustic metamaterial designs for bandgap targets \@slowromancapvi@ and \@slowromancapvii@ (marked as gray-shaded frequency regions). Each row shows five designs with the highest likelihood of satisfying a target. All the radii ($r_{\text{strut}}$, $r_{\text{center}}$, and $r_{\text{corner}}$) have a unit of $\mu$m.
Figure S5: Geometries and corresponding dispersion relations of generated acoustic metamaterial designs for bandgap targets \@slowromancapviii@ and \@slowromancapix@ (marked as gray-shaded frequency regions). Each row shows five designs with the highest likelihood of satisfying a target. All the radii ($r_{\text{strut}}$, $r_{\text{center}}$, and $r_{\text{corner}}$) have a unit of $\mu$m.
Figure S6: Geometries and corresponding dispersion relations of generated acoustic metamaterial designs for bandgap target \@slowromancapx@ (marked as gray-shaded frequency regions). Each row shows five designs with the highest likelihood of satisfying a target. All the radii ($r_{\text{strut}}$, $r_{\text{center}}$, and $r_{\text{corner}}$) have a unit of $\mu$m.
@@
Figure S7: Designs and corresponding absorbance spectra of generated optical metasurface designs for design targets \@slowromancapii@-\@slowromancapiv@ (marked as gray-shaded wavelength regions). Each row shows five designs with the highest likelihood of satisfying a target. All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm.